COMPUTATIONAL MOLECULAR SPECTROSCOPY: THE EQUILIBRIUM BOND LENGTH OF X̃^{3}Φ CoH

TSUNEO HIRANO and \underline{UMPEI NAGASHIMA}, Research Institute for Computational Sciences, National Institute of Advanced Industrial Science and Technology, 1-1-1 Umezono, Tsukuba, Ibaraki 305-8568, Japan; MUTSUMI TOMONARI, Fundamental and environmental Research Laboratories, NEC Corporation, Miyuki-gaoka 34, Tsukuba 305-8501, Japan; KIYOSHI TANAKA, Division of Chemistry, Graduate School of Science, Hokkaido University, Kita 10, Nishi 8, Kita-ku, Sapporo, 060-0810 Japan; PER JENSEN, Theoretische Chemie, Bergische Universität, D-42097 Wuppertal, Germany.

Two distinctly different values of r_{e}(Co-H) have been reported for the CoH molecule: One is 1.5138435(80)^{a} Å from the Beaton and Brown group (BB), determined by far-infrared laser magnetic resonance spectroscopy, and the other is 1.531291(8)^{b} Å or 1.532664(16)^{c} Å from the Ram and Bernath group (RB) determined by near-infrared emission spectroscopy. The BB group analyzed spectra at the Hund’s case (a) limit and the RB group at the Hund’s case (c). The rotational constants determined by the two groups are similar in magnitude: B_{Ω=4} and B_{Ω=3} by BB are 7.13680(16) and 7.27450(13) cm^{−1}, respectively, and B_{Ω=4} and B_{Ω=3} by RB in 2006 are 7.136591(160) and 7.27614(21) cm^{−1}, respectively. The difference in r_{e}(Co-H) is caused by the fact that the BB group apply the spin-orbit correction B_{Ω} = B_{0} + 2B_{0}^{2} Σ/(A_{SO}Λ) to the rotational constant, whereas the RB group do not. It should be noted that although the ^{3}Φ_{Ω=3} state is perturbed, we can derive almost the same r_{e}(Co-H) values of 1.5179 Å and 1.5180 Å from the B_{Ω=3} values reported from BB and RB groups, respectively. Here, spin-orbit correction is not needed since Σ = 0. Our best ab initio value of r_{e} is 1.507 Å (J. Chem. Phys., in press) and 1.509 Å (unpublished) at the level of MRCPA(4)/STO using the ALCHEMY II program. Since BB used the spin-orbit coupling constant A_{SO} = −221.5 cm^{−1} determined from the experimentally observed spin-splitting, the unperturbed A_{SO} becomes larger than theirs, and hence the r_{e} value reported by BB group based on B_{Ω=4} should be a little shorter than 1.5138 Å, which coincides with our ab initio predicted value.

aS. P. Beaton, K. M. Evenson, and J. M. Brown, J. Mol. Spectroscopy, 164, 395-415 (1994).

bR. S. Ram, P. F. Bernath, and S. P. Davis, J. Mol. Spectroscopy, 175, 1-6 (1996).

cI.E. Gordon, R.L. LeRoy, and P.F. Bernath, J. Mol. Spectroscopy, 237, 11-18 (2006).